OpenStudy (anonymous):

Use the substitution method to find the solution(s) to the system of equations. Show all work for full credit. y = –x^2 y = x

OpenStudy (anonymous):

x = -x^2 -x = x^2 sqrt(-x) = x the only way this could work is if x = 0

OpenStudy (anonymous):

incorrect!

OpenStudy (anonymous):

0 or -1.

OpenStudy (anonymous):

Incorrect?

OpenStudy (anonymous):

no, no, not you, meeeerah

OpenStudy (anonymous):

i don't get it, i wrote it few minutes ago, where it dissapeared? :/

OpenStudy (anonymous):

I didn't delete anything lol. & What is incorrrect?

OpenStudy (anonymous):

x = -x^2 -x = x^2 sqrt(-x) = x the only way this could work is if x = 0 - this is incorrect...i didn't write it to you

OpenStudy (anonymous):

\[-x^2=x\] \[x^2+x=0\] \[x(x+1)=0\] \[x=0\] or \[x=-1\]

OpenStudy (anonymous):

Thank you, SATELLITE73.

OpenStudy (anonymous):

that is the worked out answer, but you can probably do this with your eyeballs. 0 obviously works, as does -1. we say "by inspection" but you may not get full credit without the details

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

I know, but the teacher needs the work, thats why in an earlier comment i said 0 or -1, I just didn't know how to get that answer.

OpenStudy (anonymous):

well if you want to write even more you can put \[x+1=0\] \[x=-1\] i skipped that step. teachers should encourage answers by inspection, especially when they are rigth

OpenStudy (anonymous):

my bad